[clas_members] [New Logentry] RC Daily Summary August 7, 2022

bueltman at jlab.org bueltman at jlab.org
Sun Aug 7 16:20:01 EDT 2022

Logentry Text:
Data taking on ND3, polarization about +0.64, without sector 3, at the beginning of day shift.

Access mid-morning to replace faulty FTOF/PCAL HV mainframe sector 3 and VME power supply for FTOF4.

Continue data taking on ND3, polarization about +0.64.
Switch HWP to OUT during swing shift.

Liquid helium bath level dropped mid-shift, ice on run valve, target group was called and made an access to remove ice and restart polarizing.
Buffer dewar pressure was attempted to be lowered (to 1250 mbar), but could not be controlled by EPICS. Reboot of target ioc did not help. 

During owl data taking on ND3, polarization increasing from +0.5 and +0.6 during shift.

Pass change for Hall C might be postponed until Tuesday.

We will have no RC meeting today.

Data taking progress:
Day 4.0 nA on ND3 (HWP IN), completed runs 16583 (28 M, no sector 3), 16585 (5 M), 16586 (22 M), 16587 (27 M)
Swing 4.0 nA on ND3 (HWP IN), completed runs 16588 (67 M)
Change to HWP OUT, completed runs 16594 (21 M)
Owl 4.0 nA on ND3 (HWP OUT), completed runs 16597 (58 M), 16598 (2 M), 16599 (68 M)


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4025449
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220807/d76a4bc2/attachment.html>

More information about the clas_members mailing list